OpenStudy (anonymous):

the function defined by D(t)=13t^2-73t gives the distance in feet that a car going approximately 50mph will ksid t seconds. Find the time it would take the care to skid 334fr.

OpenStudy (bahrom7893):

You're given the function that gives you the DISTANCE in feet that a car would skid in t amount of TIME Find the TIME it takes for the car to skid a DISTANCE of 334 feet.

OpenStudy (bahrom7893):

DISTANCE = D(t) = 13*t^2 - 73t = 334 TIME = t - ? Solve: 13t^2 - 73t = 334 for t, that will be your time.